Robin Briggs, a wealthy private investor, had been

approached by Union Finance Company on the

previous day. It seemed that Union Finance was

interested in loaning money to one of its larger clients,

but the client’s demands were such that Union could

not manage the whole thing. Specifically, the client

wanted to obtain a loan for $385,000, offering to repay

Union Finance $100,000 per year over 7 years.

Union Finance made Briggs the following

proposition. Since it was bringing Briggs business, its

directors argued, they felt that it was only fair for

Briggs to put up a proportionately larger share of the

money. If Briggs would put up 60% of the money

($231,000), then Union would put up the remaining

40% ($154,000). They would split the payments

evenly, each getting $50,000 at the end of each year for

the next 7 years.

a. Union Finance can usually earn 18% on its

money. Using this interest rate, what is the net

present value of the client’s original offer to

Union?

b. Robin Briggs does not have access to the same

investments as Union. In fact, the best available

alternative is to invest in a security earning

10% over the next 7 years. Using this interest

rate, what is Briggs’s net present value of the offer made by Union? Should Briggs accept the offer?

c. What is the net present value of the deal to

Union if Briggs participates as proposed?

d. The title of this case study is “The Value of

Patience.” Which of these two investors is more

patient? Why? How is this difference exploited

by them in coming to an agreement?
Answer #1

a.Net Present Value = Present Value of Cash Inflows – Present Value of Cash Outflows

NPV of Original Offer = 100,000*PVAF(18%, 7 years) – 385,000

= 100,000*3.812 – 385,000

= $(3,800)

b.NPV = 50,000*PVAF(10%, 7 years) – 231,000

= 50,000*4.868 – 231,000

= $12,400

Since NPV is positive, Briggs should accept the offer.

c.NPV = 50,000*PVAF(18%, 7 years)- 154,000

= 50,000*3.812 – 154,000

= $36,600

d.Briggs is more patient since he is willing to contribute more because he knows that in the long term it will benefit him.